# Diffusion2: Dynamic 3D Content Generation via Score Composition of Orthogonal Diffusion Models

### v2

@@ 3.1 Image matrix generation
Since $p\left({\color[rgb]{0,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{0,.5,.5}\hat{\mathcal{I}}_{-i,-j}}|{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\hat{\mathcal{I}}_{-i,j}},{\color[rgb]{1,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,1}\pgfsys@color@cmyk@stroke{0}{1}{0}{0}\pgfsys@color@cmyk@fill{0}{1}{0}{0}\hat{\mathcal{I}}_{i,-j}}\right)$ does not contain $I_{i,j}$, its derivative with respect to $I_{i,j}$ is zero. Then combined with equation (8) and equation (9), taking the derivative of $\log p\left(\hat{\mathcal{I}}\right)$ with respect to $\mathrm{x}$, we achieve
@@
$=\nabla_{\mathrm{x}}\log p\left(\hat{I}_{i,j},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\hat{\mathcal{I}}_{-i,j}},{\color[rgb]{1,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,1}\pgfsys@color@cmyk@stroke{0}{1}{0}{0}\pgfsys@color@cmyk@fill{0}{1}{0}{0}\hat{\mathcal{I}}_{i,-j}}\right).$ (11)
@@
$=\nabla_{\mathrm{x}}\log p\left({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\hat{\mathcal{I}}_{-i,j}}|\hat{I}_{i,j}\right)p\left({\color[rgb]{1,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,1}\pgfsys@color@cmyk@stroke{0}{1}{0}{0}\pgfsys@color@cmyk@fill{0}{1}{0}{0}\hat{\mathcal{I}}_{i,-j}}|\hat{I}_{i,j}\right)p\left(\hat{I}_{i,j}\right)$ (13)
$=\nabla_{\mathrm{x}}\log\frac{p\left(\hat{\mathcal{I}}_{\{1:V\},j}\right)p\left(\hat{\mathcal{I}}_{i,\{1:F\}}\right)}{p\left(\hat{I}_{i,j}\right)}$ (14)
$=\nabla_{\mathrm{x}}\log p\left(\hat{\mathcal{I}}_{\{1:V\},j}\right)+\nabla_{\mathrm{x}}\log p\left(\hat{\mathcal{I}}_{i,\{1:F\}}\right)-\nabla_{\mathrm{x}}\log p\left(\hat{I}_{i,j}\right).$ (15)
Here $\nabla_{\mathrm{x}}\log p\left(\hat{\mathcal{I}}_{i,\{1:F\}}\right)$ and $\nabla_{\mathrm{x}}\log p\left(\hat{\mathcal{I}}_{\{1:V\},j}\right)$ are just the score functions of the video diffusion model and the multi-view diffusion model respectively . And we use the convex combination of them to estimate $\nabla_{\mathrm{x}}\log p\left(\hat{I}_{i,j}\right)$ as:
@@
In practice, we employ a logistic schedule to adjust the change of $s$ with denoising step. Given the current denoising step $i$ and the number of total steps $N$, we set $s=1-\frac{1}{1+e^{k(i/N-s_{0})}}$. This function has a sigmoidal curve, which is relatively flat at the extremes away from middle $s_{0}$ and changes sharply near it, with the derivative controlled by $k$. This schedule decouples the generation of consistent geometry and temporally smooth appearance to some extent.
Sampling in latent space
@@
For convenience, the above theorem assumes that the sample object is the image in the original RGB space. However, modern high-resolution diffusion models typically generate images encoded into a latent space by VQVAE van2017neural . The legitimacy of the aforementioned derivation requires that the multi-view generator and the video generator share the same latent space. Although this requirement is not met in the most current instantiation of them, we believe that this condition will be increasingly satisfied by more multi-view generation models in the future. Because as pointed out by SVD blattmann2023stable ; chen2024v3d ; videoworldsimulators2024 ; li2024sora , video generation models trained on large-scale video datasets have learned a strong stereo knowledge, thus can provide a better pretraining for fine-tuning multi-view diffusion models than those trained solely on image data. And the latent encoder is usually frozen during fine-tuning on the multi-view images.
Generation with various input conditions
@@
Note that the formulation described above is based on unconditional generation. However, we are more interested in controllable generation in practice. Then we will extend the aforementioned process to conditional generation. Consider the augmented matrix $\mathcal{I}_{\text{aug}}$ defined by
$\mathcal{I}_{\text{aug}}=\begin{bmatrix}I_{0,0}&\mathcal{I}_{0,\{1:V\}}\\ \mathcal{I}_{\{1:F\},0}&\mathcal{I}\end{bmatrix},$ (17)
where $I_{0,0}$ are the input image, and $\mathcal{I}_{0,\{1:V\}},\mathcal{I}_{\{1:F\},0}$ is the first row and column of $\mathcal{I}_{\text{aug}}$ we need to first create as the condition for the subsequent generation of full matrix. Then we will demonstrate how we obtain them from various conditions. For convenience, we denote $\mathbf{V}$ as the multi-view diffusion model and $\mathbf{F}$ as the video diffusion model.
  • Single image. Given $I_{0,0}$ as input, we use $\mathbf{V}$ to generate $\mathcal{I}_{0,\{1:V\}}$ which dictates the geometry of the generated 4D assets and use $\mathbf{F}$ to generate $\mathcal{I}_{\{1:F\},0}$ which endows the static image with dynamics.
@@
  • Single-view video. Given $\mathcal{I}_{\{0:F\},0}$ as input, we use the last frame $I_{0,0}$ as the condition of $\mathbf{V}$ to generate $\mathcal{I}_{0,\{1:V\}}$.
@@
  • Static 3D model. Similarly, Given $\mathcal{I}_{0,\{0:V\}}$ as input, we use the front view $I_{0,0}$ as the condition of $\mathbf{F}$ to generate $\mathcal{I}_{\{1:F\},0}$.
@@
Assumption 3.1 ensures the safety of independently generating the geometry $\mathcal{I}_{0,\{1:V\}}$ and the motion $\mathcal{I}_{\{1:F\},0}$. Additionally, there is no computational or data dependency between these two generation processes, allowing their total time cost to be reduced to a single reverse diffusion process. Then we will denoise the rest part of $\mathcal{I}_{\text{aug}}$ from pure Gaussian noise. In each step, we run the score estimators for each row and column conditioned on the $\mathcal{I}_{0,\{1:V\}}$ and $\mathcal{I}_{\{1:F\},0}$, and combine their results as in Theorem 3.1 to update the noisy latent. Since the score estimation for each row and column can also be parallelized, the time cost can be decreased to running a single diffusion step. Therefore, with sufficient GPU memory, the total time spent on the process illustrated in Figure 2(ii) remains the same as that for generating a single video.
@@ 3.2 Robust reconstruction
4D representation
Given generated synchronized multi-view videos conditioned on any type of prompts, there are numerous methods that can be employed to reconstruct 4D assets. Among the numerous candidates, we adopt the 4D Gaussian Splatting due to its superior fitting capabilities and efficient optimization given the dense multi-view supervision.
Optimization
Although the images generated in the first stage already have intuitively satisfactory spatiotemporal consistency, the performance limitation of the foundational multi-view generation components still makes it difficult to achieve precise pixel-level matching across different views and frames. Therefore, we follow chen2024v3d to optimize the combination of perception loss $\mathcal{L}_{lpips}$ zhang2018unreasonable and D-SSIM $\mathcal{L}_{ssim}$ wang2004image while ignoring L1 loss. In addition, we weight each term with the confidence score, then the total objective is defined as $\mathcal{L}_{total}=\lambda_{lpips}\mathcal{C}_{lpips}\mathcal{L}_{lpips}+\lambda_{ssim}\mathcal{C}_{ssim}\mathcal{L}_{ssim}$, where $\mathcal{C}_{ssim}$ is just the ssim between ground truth and rendered images and $\mathcal{C}_{lpips}$ is defined as $1-\mathcal{L}_{lpips}$.
@@ 4.1 Implementation details
In the first stage, we use Stable Video Diffusion blattmann2023stable as our foundation video diffusion model, predicting 25 frames each time according to the image prompt. SV3Dp voleti2024sv3d is chosen as the foundation multi-view diffusion model. For simplicity, we only generate orbital videos that have 21 uniformly spaced azimuths and fixed elevation with manual filter of side view as these views typically contain thin structures that pose challenges for the video generation model and subsequent reconstruction processes. By default, we set the number of sampling steps to 50 for both generative models. In the reconstruction stage, we optimized 4D Gaussian Splatting for 5k iterations without bells and whistles. The image size is set to (576 $\times$ 576) in both stages.
@@
4.2 4D generation from single image
@@
Figure 3 (a), we show the results generated by the proposed method and provide the comparison with other alternatives. It can be observed that our concise and elegant pipeline is capable of generating 4D assets of comparable quality to those produced by state-of-the-art SDS-based methods with sophisticated multi-stage optimization. Furthermore, the parallel evaluation of a large 2D generative model can provide a potential efficiency advantage for our method. We also conducted a user study (See appendix for details), the results of which are reported in Table 1. It suggests that our method garnered the highest human preference in the multi-view consistency, detail, and overall model quality.
4.3 4D generation from single view video
@@
Generating 4D dynamic objects from fixed-view video is a practical task first introduced in jiang2023consistent4d . Compared to generation from a single image, this task additionally constrains the object’s motion. Our proposed framework can be easily adapted to deal with this task as detailed in Section 3.1. We perform both quantitative and qualitative comparisons with other counterparts under this setting. The qualitative result is shown in Figure 3 (b). It can be seen that our method slightly reduces the over-saturated appearance. For qualitative evaluation, we report the CLIP-similarity between generated views and ground truth images to indicate overall semantic consistency and the recognizable of generated images. The quantitative metrics also support our superiority.
4.4 4D generation from static 3D content
Figure 4: Synthesized images from static 3D models.
Naturally, Diffusion2 can also animate static 3D models into dynamic objects as described in Section 3.1, which has substantial practical implications. From Figure 4 we can observe that our method is capable of endowing 3D models with diverse and realistic dynamics while maintaining satisfactory temporal and geometrical consistency.
@@ 4.5 Ablation studies
Figure 5: Ablation studies. (a) The parameter controlling the the logistic schedule. (b) Different type of scale schedule. Best viewed with zoom-in.
Since previous methods have explored dynamic content reconstruction well, we mainly focus on the ablation of key design choices in the stage of image matrix generation. In Figure 5 (a), we adjusted two parameters that control the curve of the logistic schedule. The results reveal that: when only using video prior, the generated images fail to ensure consistency in perspective, geometry, and detail with other views. For example, in the second row and second column of Figure 5 (a), the dress still evenly sags down both sides of the body, unlike in the reference view where the dress flutters to the left side of the body. This indicates that without the guidance of a geometry prior, the dynamics of each view will be totally independent of each other. Another extreme is when we set $s$ to 1, that is, using only multi-view prior, where it can no longer guarantee the consistency of details between different frames of invisible views. The comparison of the last column of the first two rows of Figure 5 (a) shows a remarkable change in the shape of the tail of the hair. So we finally adopt a compromised option to set $s$ to 0.5, i.e., starting to drastically reduce the weight of the multi-view score at halfway of the denoising process. In most cases, this choice can achieve smooth temporal transitions as well as geometrical consistency. In addition, the decreasing speed of $s$ may also affect the quality of generated images, which is controlled by another parameter $k$. Therefore, we test the image generation under different $k$ given $s=0.5$. It can be seen that when $k$ is relatively small, the generated images may exhibit ghosting effects. On the other hand, when $k$ is too large, it may weaken the temporal smoothing effect brought by the video diffusion model in the early stages of denoising. Consequently, we adopt $s=0.5,k=20$ as the default setting. Finally, we also examined the impact of different noise schedules, as shown in Figure 5 (b). It can be observed that all other choices would lead to ghosting artifacts.

### v3

@@ 3.1 Image matrix generation
Since $p\left({\color[rgb]{0,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{0,.5,.5}\hat{\mathcal{I}}_{-i,-j}}|{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\hat{\mathcal{I}}_{-i,j}},{\color[rgb]{1,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,1}\pgfsys@color@cmyk@stroke{0}{1}{0}{0}\pgfsys@color@cmyk@fill{0}{1}{0}{0}\hat{\mathcal{I}}_{i,-j}}\right)$ does not contain $\mathrm{x}$ (=$I_{i,j}$), its derivative with respect to $\mathrm{x}$ is zero thus this term does not contribute to the score of $I_{i,j}$. Then combined with equation (6) and equation (7), taking the derivative of $\log p\left(\hat{\mathcal{I}}\right)$ with respect to $\mathrm{x}$, we achieve
@@
$=\nabla_{\mathrm{x}}\log p\left(\hat{I}_{i,j},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\hat{\mathcal{I}}_{-i,j}},{\color[rgb]{1,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,1}\pgfsys@color@cmyk@stroke{0}{1}{0}{0}\pgfsys@color@cmyk@fill{0}{1}{0}{0}\hat{\mathcal{I}}_{i,-j}}\right).$
@@
$=\nabla_{\mathrm{x}}\log p\left({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\hat{\mathcal{I}}_{-i,j}}|\hat{I}_{i,j}\right)p\left({\color[rgb]{1,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,1}\pgfsys@color@cmyk@stroke{0}{1}{0}{0}\pgfsys@color@cmyk@fill{0}{1}{0}{0}\hat{\mathcal{I}}_{i,-j}}|\hat{I}_{i,j}\right)p\left(\hat{I}_{i,j}\right)$
$=\nabla_{\mathrm{x}}\log\frac{p\left(\hat{\mathcal{I}}_{\{1:V\},j}\right)p\left(\hat{\mathcal{I}}_{i,\{1:F\}}\right)}{p\left(\hat{I}_{i,j}\right)}$
$=\nabla_{\mathrm{x}}\log p\left(\hat{\mathcal{I}}_{\{1:V\},j}\right)+\nabla_{\mathrm{x}}\log p\left(\hat{\mathcal{I}}_{i,\{1:F\}}\right)-\nabla_{\mathrm{x}}\log p\left(\hat{I}_{i,j}\right).\qed$
Here $\nabla_{\mathrm{x}}\log p\left(\hat{\mathcal{I}}_{i,\{1:F\}}\right)$ and $\nabla_{\mathrm{x}}\log p\left(\hat{\mathcal{I}}_{\{1:V\},j}\right)$ are just the score functions of the video diffusion model and the multi-view diffusion model respectively. Since the $\nabla_{\mathrm{x}}\log p\left(\hat{I}_{i,j}\right)$ is intractable in the conditional generation, we use the convex combination of $\nabla_{\mathrm{x}}\log p\left(\hat{\mathcal{I}}_{i,\{1:F\}}\right)$ and $\nabla_{\mathrm{x}}\log p\left(\hat{\mathcal{I}}_{\{1:V\},j}\right)$ to replace it as:
@@
In practice, we employ a logistic schedule to adjust $s$ at each denoising step. Given the current denoising step $i$ and the number of total steps $N$, we set $s=1-\frac{1}{1+e^{k(i/N-s_{0})}}$. This function has a sigmoidal curve, which is relatively flat at the extremes away from middle $s_{0}$ and changes sharply near it, with the derivative controlled by $k$. This schedule decouples the generation of consistent geometry and temporally smooth appearance to some extent.
@@
Sampling in latent space   For convenience, the above theorem assumes that we sample images in RGB space. However, modern high-resolution diffusion models typically generate images in a latent space encoded by pretrained auto-encoder kingma2014auto ; van2017neural . The legitimacy of the aforementioned derivation requires that the multi-view generator and the video generator share the same latent space. This requirement is not met by most existing instantiation of them. However, as pointed out by SVD blattmann2023stable , video generation models trained on large-scale video datasets have learned a strong stereo knowledge, thus providing a better pretraining for multi-view diffusion models than those trained solely on image data. Therefore, we believe that the requirement will be increasingly satisfied in the future.
@@
Generation with various input conditions   Note that the formulation described above is based on unconditional generation. However, we are more interested in controllable generation in practice. Then we will discuss on how to extend the above process to conditional generation. Consider the augmented matrix $\mathcal{I}_{\text{aug}}$ defined by
$\mathcal{I}_{\text{aug}}=\begin{bmatrix}I_{0,0}&\mathcal{I}_{0,\{1:V\}}\\ \mathcal{I}_{\{1:F\},0}&\mathcal{I}\end{bmatrix}.$ (11)
Compared to the aforementioned unconditional generation, the augmented matrix $\mathcal{I}_{\text{aug}}$ additionally includes the optional input image $I_{0,0}$, and $\mathcal{I}_{0,\{1:V\}},\mathcal{I}_{\{1:F\},0}$ that represent the reference geometry and dynamic of target object. To incorporate them into the generation of $\mathcal{I}$, we condition the score estimator for denoising each row (or column) in equations (5) and (10) on the first element of the same row (or column) in $\mathcal{I}_{\text{aug}}$. This formulation enables us to handle various 4D generation tasks depending on different forms of conditions specified by users as below:
  • Image-to-4D. Given $I_{0,0}$ as input, both $\mathcal{I}_{0,\{1:V\}}$ and $\mathcal{I}_{\{1:F\},0}$ are generated by diffusion models.
@@
  • Video-to-4D. Given $\mathcal{I}_{\{0:F\},0}$ as input, we use the last frame $I_{0,0}$ as the condition of multi-view diffusion model to generate $\mathcal{I}_{0,\{1:V\}}$.
@@
  • 3D-to-4D. Similarly, Given $\mathcal{I}_{0,\{0:V\}}$ as input, we use the front view $I_{0,0}$ as the condition of video diffusion model to generate $\mathcal{I}_{\{1:F\},0}$.
  • Text-to-4D. The above inputs including single image $I_{0,0}$, single-view video $\mathcal{I}_{\{0:F\},0}$ and static multi-view images $\mathcal{I}_{0,\{0:V\}}$ can all be generated guided by text.
Once $\mathcal{I}_{0,\{1:V\}},\mathcal{I}_{\{1:F\},0}$ are obtained, we can sample the full matrix $\mathcal{I}$ using equations (5) and (10).
@@
Parallel denoising   Assumption 3.1 ensures the safety of independently generating the geometry $\mathcal{I}_{0,\{1:V\}}$ and the motion $\mathcal{I}_{\{1:F\},0}$. Since these two generation processes have no computational or data dependency, their total time cost could be reduced to a single reverse diffusion process. Additionally, when we denoise the rest part of $\mathcal{I}_{\text{aug}}$, i.e., $\mathcal{I}$, the score estimation for each row and column can also be parallelized in each denoising step. Therefore, with sufficient GPU memory, the total time spent on the process illustrated in Figure 2 (ii) remains the same as that for generating a single video.
@@ 3.2 Robust reconstruction
4D representation   Given generated synchronized multi-view videos from any type of condition, there are numerous methods that can be employed to reconstruct complete 4D assets. Among these candidates, we adopt the 4D Gaussian Splatting yang2023gs4d due to its superior fitting capabilities and efficient optimization given the dense multi-view supervision.
Optimization   Although the images produced in the first stage already exhibit intuitively satisfactory spatiotemporal consistency, the performance limitation of the foundational diffusion model still makes it difficult to achieve precise pixel-level matching across different views and frames. Therefore, we follow chen2024v3d to optimize the combination of perception loss $\mathcal{L}_{lpips}$ zhang2018unreasonable and D-SSIM $\mathcal{L}_{ssim}$ wang2004image while discarding L1 loss. In addition, we weight each term with the confidence score as pan2024fast , then the total objective is defined as $\mathcal{L}_{total}=\lambda_{lpips}\mathcal{C}_{lpips}\mathcal{L}_{lpips}+\lambda_{ssim}\mathcal{C}_{ssim}\mathcal{L}_{ssim}$, where $\mathcal{C}_{ssim}$ is just the SSIM between ground truth and rendered images and $\mathcal{C}_{lpips}$ is defined as $1-\mathcal{L}_{lpips}$.
@@ 4.1 Implementation details
[…]
@@
4.2 Main results
4D generation from reference image
@@ 4.2 Main results
In Figure 3 (a), we present the results generated by our method and compare them with other alternatives. It can be observed that our simple and elegant pipeline is capable of generating 4D assets with comparable quality to those produced by SDS-based methods using sophisticated multi-stage optimization. Furthermore, the parallel denoising characteristic can provide an efficiency advantage for our method over those need recurrent evaluation of diffusion UNet. We also conducted a user study in Table 1. It suggests that our method garnered the highest preference in the multi-view consistency, detail, and overall model quality.
4D generation from single view video   Generating 4D dynamic objects from single-view video is a practical task first introduced in jiang2023consistent4d . Compared to image-to-4D generation, this task places additional constraints on the object’s motion. Our proposed framework can be easily adapted to deal with this task as detailed in Section 3.1. We perform both quantitative and qualitative comparisons with other counterparts in this setting. The qualitative result is shown in Figure 3 (b). It can be seen that our method slightly reduces the over-saturated appearance led by SDS while achieving more clear details. For quantitative evaluation, we report the CLIP-similarity between generated views and ground truth images to indicate overall semantic consistency and the recognizability of generated images. As shown in Table 2, our method can achieve better CLIP similarity than other alternatives.
4D generation from static 3D content
Figure 4: Synthesized images from static 3D models.
@@
Naturally, Diffusion2 can also animate static 3D models as described in Section 3.1, which has substantial practical implications. From Figure 4 we can observe that our method is capable of endowing 3D models with diverse and realistic dynamics while maintaining satisfactory temporal and geometrical consistency. More results including text-to-4D can be found in Section A.5.
@@ 4.3 Ablation studies
Figure 5: Ablation studies. (a) & (b) The tuning of parameters $s_{0}$ and $k$ controlling the logistic schedule. (c) Different types of scale schedule. Best viewed with zoom-in.
Since previous methods have explored dynamic content reconstruction well, we mainly focus on the ablation of key design choices in the stage of image matrix generation. Recall that in equation (10), we use $s$ to balance the contribution of two foundation models. And we choose a logistic function $s=1-\frac{1}{1+e^{k(i/N-s_{0})}}$ as the schedule of $s$, which has two hyper-parameters $s_{0}$ and $k$ controlling the shape of the function. Then we tune $s_{0}$ and $k$ to study their impact on generation.
Impact of $s_{0}$   In Figure 5 (a), we tune $s_{0}$ with $k$ fixed to $20$. We first consider two extreme cases. In the first case, we set $s_{0}=0$, implying that $s<0.5$ all the time and SVD has dominant contribution in equation (5). As shown in “$s_{0}=0,k=20$” of Figure 5 (a), the dress still evenly sags down both sides of the body, unlike in the reference view where the dress flutters to the left side of the body. This indicates that without the sufficient guidance of geometry prior, the dynamics of each view will be independent of each other. Another extreme is when we set $s_{0}=1$, where SV3D has dominant contribution. In this case, it can no longer guarantee the consistency of details between different frames of invisible views. The comparison of the images of “$s_{0}=1,k=20$” shows a remarkable change in the shape of the tail of the hair. By further telescoping the results with different $s_{0}$, we finally adopt a compromised option to set $s_{0}$ to $0.5$, i.e., starting to drastically reduce the weight of the multi-view score at halfway of the denoising process. In most cases, this choice can achieve a satisfactory trade-off between smooth temporal transitions and geometrical consistency.
Impact of $k$   In addition, the decreasing speed of $s_{0}$ may also affect the quality of generated images, which is controlled by another parameter $k$. Therefore, we test the image generation under different $k$ given $s_{0}=0.5$ in Figure 5 (b). It can be seen that when $k$ is relatively small, the generated images may exhibit ghosting effects. On the other hand, when $k$ is too large, it may weaken the temporal smoothing effect brought by the video diffusion model in the early stage of denoising process. Consequently, we adopt $s=0.5,k=20$ as the default setting.
Impact of logistic schedule   Finally, we also examined the impact of different schedule types, as shown in Figure 5 (c). We compare with two counterparts: “Constant” ($s\equiv 0.5$) and “Linear” ($s=\frac{i}{N}$). It can be observed that other alternates would lead to ghosting artifacts.